# Mousse: Rectifying the Geometry of Muon with Curvature-Aware Preconditioning

### 10 Mar 2026 v1

@@ 1 Introduction
Optimizing large language models requires navigating a high-dimensional and notoriously ill-conditioned loss landscape [sagun2016eigenvalues]. While adaptive optimization has firmly established the bedrock of deep learning — spanning from the ubiquitous coordinate-wise AdamW [loshchilov2017decoupled] to advanced structural preconditioners like Shampoo [gupta2018shampoo] and SOAP [vyas2024soap] — these methods primarily focus on refining the gradient direction within Euclidean geometry. Recently, a new paradigm of spectral optimization has gained significant traction in both academia and industry. Notably, the Muon optimizer [jordan2024muon] has emerged as a compelling alternative, constraining updates to the Stiefel manifold by functionally enforcing global spectral regularization via Newton-Schulz iterations. Leading AI Laboratories, including Moonshot-AI [team2025kimi], DeepSeek-AI [cheng2026conditional] and ZhipuAI [zeng2025glm], have integrated Muon optimizer in their pre-training pipelines to capitalize on its superior scale-invariance and convergence efficiency at scale.
@@ 2 Related Works
3 Preliminaries
In this section, we establish the mathematical foundation of our work by framing optimization as a geometric process connecting dual spaces. We denote the parameter space as $\mathcal{W}$ and the gradient space as $\mathcal{G}\cong\mathcal{T}_{W}^{*}\mathcal{W}$. The gradient $G=\nabla_{W}\mathcal{L}\in\mathcal{G}$ represents a linear functional on the tangent space $\mathcal{T}_{W}\mathcal{W}$, and the role of an optimizer is to define a mapping $p:\mathcal{G}\rightarrow\mathcal{T}_{W}\mathcal{W}$ that translates this functional into a concrete update step $\Delta W$. We use $\text{vec}(\cdot)$ to denote the vectorization operator, which stacks the columns of a matrix into a single column vector.
@@ 3.1 Muon as Spectral Steepest Descent
$\Delta W_{\text{Muon}}=-\text{msign}(G)=-UV^{T}$ (3)
@@
where $G=U\Sigma V^{T}$ is the singular value decomposition (SVD) of the gradient. This projection is equivalent to restricting the update onto the Stiefel manifold, ensuring $\Delta W^{T}\Delta W=I$. It is worth noting that the spectral norm provides a looser constraint than the Frobenius norm used in SGD, as $\|\cdot\|_{op}\leq\|\cdot\|_{F}$ holds true for any matrix. This implies that Muon permits updates with significantly larger total energy (under Frobenius norm) compared to SGD, while still strictly bounding the maximum step size along any single eigen-direction. This capacity for more aggressive updates without compromising directional stability is a key factor in its superior convergence.
@@ 3.2 Mousse: The Geometric Framework
To rectify this, Mousse reformulates the optimization problem by imposing the spectral constraint with regards to the anisotropic curvature:
@@
It can be regarded as the inner product of $\Delta W$ under the whitened basis, giving $\text{vec}(\tilde{U})=H^{\frac{1}{2}}\text{vec}(U)$. So a geometric-aware Muon is equivalent to solving the following minimizing problem:
@@
Theoretically, different approximations of $H$ yield different algorithms. Mousse adopts Shampoo’s layerwise Kronecker-Factored approximation of Hessian as the preconditioner (Approximated as $H\approx(R\otimes L)^{\frac{1}{2}}$), which has been proved to have a good performance and comparable computation costs. $L$ and $R$ capture the correlations between gradients along the rows and columns respectively, accumulated via exponential moving averages:
@@
We can directly give the solution comparable to the Muon:
@@
6:Spectral Decomposition
7:if $t\pmod{T}=0$ then $\triangleright$ Execute every $T$ steps
@@
14:Spectral Constraint
[…]
17:Unwhitening
[…]
21:Parameter Update
@@ 3.3 Comparison with Muon and SOAP
Figure 3: Validation Loss comparison on FineWeb (20B tokens). We report the final validation loss of Mousse against AdamW, Muon, and SOAP across varying peak learning rates for model sizes ranging from 160M to 800M. Mousse consistently achieves the lowest validation loss across all model scales, demonstrating superior performance and scalability.
@@ 4.1 Settings
We conducted our experiments within a modified version of Dion [ahn2025dion] library, a solid implementation of distributed muon training. For all experiments, we utilized a standard decoder-only GPT-2 architecture with several modern modifications (Appendix C.1). To ensure stable training dynamics across different scales and to facilitate hyperparameter transfer, we adopted the spectral scaling condition [yang2022tensor] for model initialization and parameter scaling. We evaluated performance across a wide spectrum of model sizes, including 160M, 240M, 480M and 800M parameters.
@@
We compared Mousse against three strong baselines representing the state-of-the-art in adaptive and spectral optimization including AdamW, Muon and Soap. To ensure a fair comparison, we conducted an extensive grid search for the peak learning rate for every optimizer across all model sizes and settings. Matrix-wise optimizers including Muon, Soap and Mousse utilize Lion [chen2023symbolic] for embedding and lmhead modules.
@@ 4.2 Main Results
As visually summarized in Figure 3, Mousse demonstrates a strict performance advantage across all model scales. Quantitatively, on the 800M parameter model, Mousse reduces the final validation loss by approximately 0.012 compared to the best Muon baseline. By integrating Shampoo-style preconditioning, Mousse successfully exploits high-curvature directions effectively, maintaining a distinct lead in the final converged loss. (See more experiment results in Appendix B.1)
@@
Another critical advantage of Mousse lies in its sample efficiency. Mousse exhibits a significantly steeper descent trajectory once the curvature statistics stabilize, and quantitatively (see Figure 4), Mousse requires approximately 12% fewer training steps to reach the final converged loss level of the Muon baseline. Figure 5 visualizes this scalability trade-off across model sizes. The dashed lines (Total Training Time) reveal that Mousse maintains a training speed nearly identical to the highly efficient Muon optimizer, with negligible overhead. In sharp contrast, SOAP suffers from significant throughput degradation. By achieving strictly lower validation loss without the computational penalty typical of curvature-corrected methods, Mousse effectively establishes a new Pareto frontier for large-scale pre-training efficiency.
@@
Despite incorporating second-order information, Mousse incurs less computational and memory overhead than traditional method like SOAP. By relying on the spectral update via Newton-Schulz iterations, Mousse intrinsically regulates step size, eliminating the need for a full-sized FP32 variance buffer required by Adam-style optimizers. Furthermore, when combined with the single-sided preconditioning technique detailed in Section 5.3, Mousse’s memory footprint is further minimized, reducing peak memory usage to approximately 88% of SOAP, achieving memory efficiency comparable to the lightweight Muon optimizer (about 1.05x vs Muon) while delivering superior convergence. Table 1 shows an overall comparison between different optimizers.
@@ 5.1 Grafting
As shown in Figure 6, without external grafting, we observe a continuous downward drift in the update RMS norm irrespective of the cosine learning rate scheduler. This diminishing updates not only impede convergence speed but also introduce unnecessary complexity to learning rate tuning. However, applying grafting effectively counteracts this trend by maintaining a stable update magnitude. Our empirical results confirm that, the ungrafted baseline exhibits performance degradation in later training stages, attributed to insufficient or unstable updates from the decaying RMS norm. In contrast, incorporating grafting consistently yields superior performance and stability.
@@ A.1 General Framework of Geometric Spectral Descent
We have mentioned in 3.1 that different approximations of $H$ yield different algorithms under the framework of geometric spectral descent. To be specific, let’s denote the preconditioner as $\tau_{H}(\cdot)$, and the target problem can be written as:

### 01 Apr 2026 v2

@@ 1 Introduction
Optimizing large language models requires navigating a high-dimensional and notoriously ill-conditioned loss landscape [sagun2016eigenvalues]. While adaptive optimization has firmly formed the backbone of deep learning — spanning from the ubiquitous coordinate-wise AdamW [loshchilov2017decoupled] to advanced structural preconditioners like Shampoo [gupta2018shampoo] and SOAP [vyas2024soap] — these methods primarily focus on refining the gradient direction within Euclidean geometry. Recently, a new paradigm of spectral optimization has gained significant traction in both academia and industry. Notably, the Muon optimizer [jordan2024muon] has emerged as a compelling alternative, constraining updates to the Stiefel manifold by functionally enforcing global spectral regularization via Newton-Schulz iterations. Leading AI laboratories, including Moonshot-AI [team2025kimi], DeepSeek-AI [cheng2026conditional] and ZhipuAI [zeng2025glm], have integrated Muon optimizer into their pre-training pipelines to capitalize on its superior scale-invariance and convergence efficiency at scale.
@@ 2 Related Works
Source code is available at https://github.com/Anti-Entrophic/Mousse.
@@
3 Methods
In this section, we present the mathematical foundation of our method and introduce its concrete formulation by framing optimization as a geometric process connecting dual spaces. We denote the parameter space as $\mathcal{W}$ and the gradient space as $\mathcal{G}\cong\mathcal{T}_{W}^{*}\mathcal{W}$. The gradient $G=\nabla_{W}\mathcal{L}\in\mathcal{G}$ represents a linear functional on the tangent space $\mathcal{T}_{W}\mathcal{W}$, and the role of an optimizer is to define a mapping $p:\mathcal{G}\rightarrow\mathcal{T}_{W}\mathcal{W}$ that translates this functional into a concrete update step $\Delta W$. We use $\text{vec}(\cdot)$ to denote the vectorization operator, which stacks the columns of a matrix into a single column vector.
@@ 3.1 Muon as Spectral Steepest Descent
$\Delta W_{\text{Muon}}=-\text{msign}(G)=-U_{\text{svd}}V_{\text{svd}}^{T}$ (3)
@@
where $G=U_{\text{svd}}\Sigma V_{\text{svd}}^{T}$ is the singular value decomposition (SVD) of the gradient. This projection is equivalent to restricting the update onto the Stiefel manifold, ensuring $\Delta W^{T}\Delta W=I$. It is worth noting that the spectral norm provides a looser constraint than the Frobenius norm used in SGD, as $\|\cdot\|_{op}\leq\|\cdot\|_{F}$ holds true for any matrix. This implies that Muon permits updates with significantly larger total energy (under Frobenius norm) compared to SGD, while still strictly bounding the maximum step size along any single eigen-direction. This capacity for more aggressive updates without compromising directional stability is a key factor in its superior convergence.
@@ 3.2 Mousse: The Geometric Framework
To rectify this, Mousse reformulates the optimization problem by imposing the spectral constraint with regards to the anisotropic curvature. Let $H$ denote a positive semi-definite approximation to the local curvature. Then the size of an update should be measured not in the raw Euclidean coordinates, but under the quadratic form induced by $H$:
@@
Equivalently, this geometry can be expressed in a whitened coordinate system. Defining
$\text{vec}(\tilde{U})=H^{\frac{1}{2}}\text{vec}(U),$ (5)
the quadratic form becomes the standard Euclidean norm in the transformed space. Mousse then imposes Muon’s spectral constraint in this whitened basis, yielding the following optimization problem:
@@
Theoretically, different approximations of $H$ yield different algorithms. Mousse adopts Shampoo’s layerwise Kronecker-factored Hessian approximation as the preconditioner, with $H\approx(R\otimes L)^{\frac{1}{2}}$, where $L$ and $R$ are estimated from the exponential moving averages of $GG^{T}$ and $G^{T}G$, respectively. This approximation has been shown to offer strong performance at comparable computational cost.
@@
The corresponding solution is analogous to that of Muon:
@@
6:if $t\pmod{T}=0$ then $\triangleright$ Execute every $T$ steps
@@
17:Parameter Update
@@ 3.3 Comparison with Muon and SOAP
Figure 3: Validation Loss comparison on FineWeb (20B tokens). We report the final validation loss of Mousse against AdamW, Muon and SOAP across varying peak learning rates for model sizes ranging from 160M to 800M. Mousse consistently achieves the lowest validation loss across all model scales, demonstrating superior performance and scalability.
@@ 4.1 Settings
We conducted our experiments within a modified version of Dion [ahn2025dion] library, a solid implementation of distributed Muon training. For all experiments, we utilized a standard decoder-only GPT-2 architecture with several modern modifications (Appendix C.1). To ensure stable training dynamics across different scales and to facilitate hyperparameter transfer, we adopted the spectral condition [yang2022tensor] for model initialization and parameter scaling. We evaluated performance across a wide spectrum of model sizes, including 160M, 240M, 480M and 800M parameters.
@@
We compared Mousse against three representative baselines in adaptive and spectral optimization including AdamW, Muon and SOAP. To ensure a fair comparison, we conducted an extensive grid search for the peak learning rate for every optimizer across all model sizes and settings. Matrix-wise optimizers including Muon, SOAP and Mousse utilize Lion [chen2023symbolic] for embedding and lmhead modules.
@@ 4.2 Main Results
As visually summarized in Figure 3, Mousse demonstrates a strict performance advantage across all model scales. Quantitatively, on the 800M parameter model, Mousse reduces the final validation loss by approximately 0.012 compared to the best Muon baseline. By integrating Shampoo-style preconditioning, Mousse exploits high-curvature directions effectively, maintaining a distinct lead in the final converged loss. (See more experiment results in Appendix B.1)
@@
Another critical advantage of Mousse lies in its sample efficiency. Mousse exhibits a significantly steeper descent trajectory once the curvature statistics stabilize. Quantitatively (see Figure 4), Mousse requires approximately 12% fewer training steps to reach the final converged loss level of the Muon baseline. Figure 5 visualizes this scalability trade-off across model sizes. The dashed lines (Total Training Time) reveal that Mousse maintains a training speed nearly identical to the highly efficient Muon optimizer, with negligible overhead. In sharp contrast, SOAP suffers from significant throughput degradation. By achieving strictly lower validation loss without the computational penalty typical of curvature-corrected methods, Mousse effectively establishes a new Pareto frontier for large-scale pre-training efficiency.
@@
Despite incorporating second-order information, Mousse incurs less computational and memory overhead than traditional methods like SOAP. By relying on the spectral update via Newton-Schulz iterations, Mousse intrinsically regulates step size, eliminating the need for a full-sized FP32 variance buffer required by Adam-style optimizers. Furthermore, when combined with the single-sided preconditioning technique detailed in Section 5.3, Mousse’s memory footprint is further reduced, bringing peak memory usage to approximately 88% of that of SOAP. This is comparable to the lightweight Muon optimizer (about 1.05x that of Muon), while still delivering superior convergence. Table 1 shows an overall comparison between different optimizers.
@@ 5.1 Grafting
As shown in Figure 6, without external grafting, we observe a continuous downward drift in the update RMS norm irrespective of the cosine learning rate scheduler. These diminishing updates not only impede convergence speed but also introduce unnecessary complexity to learning rate tuning. However, applying grafting effectively counteracts this trend by maintaining a stable update magnitude. Our empirical results confirm that the ungrafted baseline exhibits performance degradation in later training stages, attributed to insufficient or unstable updates from the decaying RMS norm. In contrast, incorporating grafting consistently yields superior performance and stability.
@@ A.1 General Framework of Geometric Spectral Descent
We have mentioned in 3.2 that different approximations of $H$ yield different algorithms under the framework of geometric spectral descent. To be specific, let’s denote the preconditioner as $\tau_{H}(\cdot)$, and the target problem can be written as: